# SpatialTree: How Spatial Abilities Branch Out in MLLMs

### v1

@@ 1 Introduction
Spatial abilities refer to the capacity to perceive, understand, reason about, and interact with 2D and 3D space, a long-standing topic in cognitive science [48, 45, 13]. In multimodal large language models (MLLMs), these abilities form the cornerstone of Spatial Intelligence (SI), yet remain challenging to study systematically due to their inherent complexity and broad scope [31, 63].
Recent research followed a task-centric trajectory. Early efforts concentrated on spatial tasks within single images [34, 50, 9], such as relative object positioning and size estimation. As capabilities matured, studies expanded into the 3D domain, tackling grounding, detection, and captioning from point clouds [67, 15, 57]. More recently, the advent of multi-view and video-capable models has further diversified the landscape [60, 52, 51, 10, 64, 20], covering a broad spectrum from spatial relevant reasoning to egocentric and dynamic object understanding.
@@
Drawing inspiration from cognitive science insight that "intelligence is a dynamic structure built through successive stages" [40], we advocate a paradigm shift: moving from fragmented tasks to a capability-centric framework. Specifically, we structure Spatial Abilities as a four level capability tree (Fig. 1):
@@
To populate this taxonomy, we first reorganize and unify data from prior works [64, 60, 20, 28, 51, 50, 67, 65, 58, 30], which primarily cover specific sub-domains within L1 to L3. To address missing abilities and enrich annotations beyond existing datasets, we construct a Spatial Engine that integrates multiple expert models [61, 54, 56, 36, 49, 27, 55, 32]. For L1 to L3, we enhance data diversity by introducing multiple QA formats for the same underlying problem and by incorporating additional sub-capabilities, including orientation estimation, localization, and affordance understanding. In contrast, L4 requires agentic interactions that are largely absent from prior benchmarks. We therefore meticulously curate L4 data across three representative embodiments: character navigation, robot grippers, and human hands. To facilitate MLLM adaptation to agentic tasks, we design an action mapping strategy that discretizes low-level actions into high-level motion primitives, forming an executable action space. In addition, human–object interaction sequences are reformatted into multi-step multiple-choice tasks through manual annotation. Based on the resulting datasets and task formulations, we establish SpatialTree-bench to evaluate a broad range of mainstream open-source and commercial MLLMs.
@@ 2 Related Work
Understanding spatial cognition has long been a central goal in cognitive science and AI. A common insight from classical theories is that spatial abilities are hierarchical, ranging from basic perception and sensorimotor interactions to higher-level reasoning and planning. Piaget [40] highlighted the developmental progression of such abilities, Tolman [48] introduced the idea of cognitive maps to represent environments for flexible navigation, and Kuipers [23, 24] formalized a hierarchical spatial representation linking local perception to global knowledge. More recent symbolic and neural approaches [45, 37] extend these insights to computational models of spatial representation, memory, and reasoning. These studies collectively motivate our SpatialTree, which organizes spatial intelligence into multi-level capabilities, bridging classical theory with systematic computational evaluation.
@@
The success of GPT-3 [6] and GPT-3.5 [38] demonstrated the potential of large language models for complex linguistic understanding and reasoning. GPT-4V [39] extends GPT-4 [1] with visual inputs, enabling single-image understanding and basic spatial reasoning. Open-sourced models such as LLaVA [29] and QwenVL [2] gradually added multi-image and video capabilities, supporting spatiotemporal reasoning. Reasoning-augmented LLMs, pioneered by OpenAI O1 [19] and DeepSeek-R1 [11], integrate chain-of-thought and reinforcement learning to enhance high-level inference. Building on these advances, GPT-4O [17] and Gemini 2.5 [8] combine perception and reasoning to support complex, agentic decision-making. Collectively, these milestones progressively enable hierarchical spatial intelligence in MLLMs, motivating structured benchmarks and evaluation frameworks across low-level perception, intermediate reasoning, and high-level agentic competence.
@@
Benchmarks for spatial abilities in MLLMs have evolved alongside the models themselves. Early efforts, such as BLINK [9], SpatialEval [50], and 3DSR-Bench [34], focused on evaluating spatial understanding tasks in single images, including distance estimation, relational question answering, and spatial captions. As MLLMs increasingly support multi-frame and video inputs, benchmarks such as VSI-Bench [60] and MMSI-Bench [64] have emerged to evaluate spatial reasoning across multiple views and dynamic scenes. To further enrich task diversity and coverage, Omnispatial [20], SITE [52], and IR3D-Bench [30] extend benchmarks to geometry puzzles, dynamic reasoning, and inverse rendering tasks. Built upon prior efforts, our SpatialTree benchmark systematically organizes spatial abilities into a hierarchical framework, providing the first thorough evaluation across different capabilities.
@@ 3.4 Agentic Competence
where $O_{t}\in\mathcal{O}$ is the current multi-modal observation, $S_{t}\in\mathcal{S}$ the internal latent state (e.g., goal, plan, or belief), $A_{t}\in\mathcal{A}$ the chosen action, and $M_{t}\in\mathcal{M}$ the updated memory representation. MLLMs are expected to output interactive actions executable across 3D environments and embodiments, such as games, simulators, and the physical world. Unlike Vision-Language Action Models (VLAs) decording the low-level control signals in robotics [18], MLLMs take the language as the only interface to link with environments like GUI Agents [41].
@@ 4.1 Data Curation
Data Annotation Framework. As shown in Fig. 3, the data engines are organized hierarchically. At the perception level, we employ a set of expert perception models, including DepthAnything3 [61], SpatialTracker [56, 55], GeoCalib [49], and OrientAnything [54], to extract intermediate perceptual representations, such as depth, correspondences, tracking results, and gravity direction. These representations are then further processed using our designed QA templates and rephrased by LLMs to construct the target QAs. For L2, we leverage 3D reconstruction pipelines (Sec. C.2) to generate BEV maps from videos, which are subsequently captioned and transformed into QAs for cognitive mapping and memory retrieval with multimodal LLMs. For other understanding tasks, such as affordance, we build on partially annotated datasets [33] and augment them by randomly constructing multiple-choice candidates through various visual prompts and by reformulating questions with more abstract descriptions. For L3, we build upon annotated reasoning-related QAs and introduce structured thinking templates, enabling an LLM-based rephraser to augment the original QAs with explicit Chain-of-Thought (CoT) reasoning. For agentic tasks, we curate Internet data covering manipulation and navigation across diverse embodiments (e.g., human hands, robotic grippers, and game characters). We further process the data with our Action-Extraction Pipeline (Sec. C.2), converting embodiment-specific actions into unified key–mouse action sequences suitable for MLLMs. In parallel, we introduce a human annotation pipeline to annotate videos with executable multi-option action sequences.
Data Resources. Our SpatialTree-Bench is constructed by systematically reorganizing numerous recent datasets (detailed in the Appendix) [60, 64, 67, 50, 65, 28, 20, 62, 51, 58, 34] to address their scattered capability coverage and over-reliance on simple questions. We first map each question to our SpatialTree framework and then enhance the evaluation protocol; for instance, complex reasoning tasks from CameraBench and MMSI-Bench are converted to a hybrid multi-option + LLM-as-a-Judge format for a finer-grained assessment. To fill the remaining capability gaps, we introduce SpatialPlus, a new dataset targeting underrepresented abilities (e.g., L1 Orientation, L1 Shape, L2 Spatial Caption) with a primary emphasis on L4 Agentic Competence. We leverage our proprietary SpatialEngine to automatically create annotations from a diverse array of video sources, including 3D reconstruction datasets, in-game footage [21], egocentric manipulation videos [16], and robotics data [22]. More implementation details are discussed in Appendix.
@@ 5.1 Models and Metrics
Benchmarked Models. We categorize the evaluated MLLMs into three groups: (1) Thinking Models, i.e., models augmented with explicit reasoning or chain-of-thought generation mechanisms (reasoning-augmented), including Gemini 2.5 Flash, Gemini 2.5 Pro [8], GLM-4.5V [14], and SeedVL1.6-Vision [12]; (2) Non-Thinking Models, which do not explicitly optimize for reasoning-style generation, such as Gemini 2.5-Pro-Nonthink, Gemini 2.5-Flash-Nonthinking, and GPT-4o [17]; and (3) Open-Source Models, including Qwen2.5-VL [3], Qwen3-VL [59], and Kimi-VL [47], representing recent community-driven multimodal advances. This diverse selection enables comprehensive comparisons across reasoning and non-reasoning paradigms, proprietary and open-source ecosystems, and model scales ranging from 32B to 72B parameters—providing a holistic overview of the current MLLM landscape. A full list of evaluated models is shown in Tab. 5.2.
@@ 5.2 Overall Performance
Gemini2.5 Flash NT 6 35.8 31.6 29.3 30.8 35.2 45.4 36.4 53.7 28.4 36.9 27.6 45.7
@@
We first present the overall performance of all benchmarked models on our proposed SpatialTree-Bench, with detailed results summarized in Tab. 5.2. In our benchmark, Gemini2.5-Pro achieves the best results (50.1) and Qwen3VL-235B get 40.0 leads the benchmark among open-source models.
@@ 6.1 Probing Cross-Ability Transfer via SFT
General Data Mixture. To construct the general visual-instruction data, we follow the VST [63] data mixing recipe and combine multimodal datasets from LLaVA-Video [66], LLaVA-NeXT-Interleave [26], and LLaVA-OneVision [25], covering single-image, multi-image, video, and 3D tasks. We then use SpatialEngine to generate ability-specific instruction data, mixed with the general data in a 1:3 ratio specifically for each. To isolate the gains from general data, we use a baseline fine-tuned only on the general data with the same token consumption.
Targeted SFT Data. For L1-Geo.Dist, we generate approximately 0.25M distance-relevant QA samples from SUNRGBD [46], Hypersim [43], and Matterport3D [7]. The training data is further augmented with visual prompts and multi-scale transformations to enhance distance reasoning. For L1-Relat.Corr, we generate matching data following VST [63], sampling 0.25M examples. Similarly, for L1-Geo.Size, we generate 0.25M samples from 3D bounding-box annotated datasets, including SUNRGBD, Hypersim and ArkitScenes [5].
@@ 6.2 Reinforcement Learning
Building on the findings from Supervised Fine-Tuning (SFT), we further investigate the potential of Reinforcement Learning with Verifiable Rewards (RLVR) to scale spatial abilities. Specifically, we employ Group Relative Policy Optimization (GRPO) [44] to align the model’s policy with the hierarchical nature of the SpatialTree. Our experiments reveal that a uniform RL strategy is insufficient, conversely a hierarchy-aware reward mechanism unlocks significant performance gains.
@@ B Evaluation Metrics Details
To assess agentic competence, models are deployed in interactive simulated environments, such as those provided by EmbodiedBench [62]. We evaluate navigation and manipulation tasks along multiple dimensions: success rate in completing the target goal, relative translation accuracy, and directional alignment. For each action step, a combined metric is computed using relative distance and cosine similarity of movement vectors, producing a normalized score in $[0,1]$. Aggregating scores over all steps yields a comprehensive measure of an agent’s ability to plan and execute actions in long-horizon, embodied tasks.
@@ C SpatialPlus: Complementary Data Annotations for SpatialTree
Figure B: Orientation Annotations. The left side is the gravity field estimated from GeoCalib [49], while the right side is from OrientAnything.
@@ C.1 Orientations (L1)
The Orientation capability, a fundamental yet under-explored area, involves estimating both gravity direction and 3D object orientation. To generate annotations, we leveraged Geocalib [49] for gravity vector estimation and OrientAnything [54] for object poses. We applied these tools to datasets suited for each task: for gravity, we annotated 500 images sampled from the diverse, drone-captured TartanAir [53] dataset; for object orientation, we utilized the object-centric Co3dv2 [42] dataset (Seen in Fig. B). For gravity, the goal is to estimate the camera’s orientation relative to the gravity vector, typically represented by the pitch and roll angles. Formally, let the gravity vector in the world frame be:

### v2

@@ 1 Introduction
Spatial abilities refer to the capacity to perceive, understand, reason about, and interact with 2D and 3D space, a long-standing topic in cognitive science [tolman1948cognitive, shepard1988mental, hegarty2010components]. In multimodal large language models (MLLMs), these abilities form the cornerstone of Spatial Intelligence (SI), yet remain challenging to study systematically due to their inherent complexity and broad scope [liu2025spatial, vst_yang2025].
Recent research followed a task-centric trajectory. Early efforts concentrated on spatial tasks within single images [ma20243dsrbench, wang2024spatial, fu2024blink], such as relative object positioning and size estimation. As capabilities matured, studies expanded into the 3D domain, tackling grounding, detection, and captioning from point clouds [zhu2024llava, hong20233d, xu2024pointllm]. More recently, the advent of multi-view and video-capable models has further diversified the landscape [yang2025thinking, wang2025site, wang2025spatialviz, gholami2025spatial, yang2025mmsi, jia2025omnispatial], covering a broad spectrum from spatial relevant reasoning to egocentric and dynamic object understanding.
@@
Drawing inspiration from cognitive science insight that "intelligence is a dynamic structure built through successive stages" [piaget2013construction], we advocate a paradigm shift: moving from fragmented tasks to a capability-centric framework. Specifically, we structure Spatial Abilities as a four level capability tree (Fig. 1):
@@
To populate this taxonomy, we first reorganize and unify data from prior works [yang2025mmsi, yang2025thinking, jia2025omnispatial, lin2025towards, wang2025spatialviz, wang2024spatial, zhu2024llava, yin2025spatial, xu2025multi, liu2025ir3d], which primarily cover specific sub-domains within L1 to L3. To address missing abilities and enrich annotations beyond existing datasets, we construct a Spatial Engine that integrates multiple expert models [yang2024depth, orient_anything, xiao2025spatialtrackerv2, mao2025spatiallm, veicht2024geocalib, li2025infinityhuman, xiao2024spatialtracker, liu2025trace]. For L1 to L3, we enhance data diversity by introducing multiple QA formats for the same underlying problem and by incorporating additional sub-capabilities, including orientation estimation, localization, and affordance understanding. In contrast, L4 requires agentic interactions that are largely absent from prior benchmarks. We therefore meticulously curate L4 data across three representative embodiments: character navigation, robot grippers, and human hands. To facilitate MLLM adaptation to agentic tasks, we design an action mapping strategy that discretizes low-level actions into high-level motion primitives, forming an executable action space. In addition, human–object interaction sequences are reformatted into multi-step multiple-choice tasks through manual annotation. Based on the resulting datasets and task formulations, we establish SpatialTree-bench to evaluate a broad range of mainstream open-source and commercial MLLMs.
@@ 2 Related Work
Understanding spatial cognition has long been a central goal in cognitive science and AI. A common insight from classical theories is that spatial abilities are hierarchical, ranging from basic perception and sensorimotor interactions to higher-level reasoning and planning. Piaget [piaget2013construction] highlighted the developmental progression of such abilities, Tolman [tolman1948cognitive] introduced the idea of cognitive maps to represent environments for flexible navigation, and Kuipers [kuipers1978modeling, kuipers2000spatial] formalized a hierarchical spatial representation linking local perception to global knowledge. More recent symbolic and neural approaches [shepard1988mental, newcombe2000making] extend these insights to computational models of spatial representation, memory, and reasoning. These studies collectively motivate our SpatialTree, which organizes spatial intelligence into multi-level capabilities, bridging classical theory with systematic computational evaluation.
@@
The success of GPT-3 [brown2020language] and GPT-3.5 [openai-gpt-3.5] demonstrated the potential of large language models for complex linguistic understanding and reasoning. GPT-4V [openai2023gpt4v] extends GPT-4 [achiam2023gpt] with visual inputs, enabling single-image understanding and basic spatial reasoning. Open-sourced models such as LLaVA [liu2023visual] and QwenVL [bai2023qwen] gradually added multi-image and video capabilities, supporting spatiotemporal reasoning. Reasoning-augmented LLMs, pioneered by OpenAI O1 [jaech2024openai] and DeepSeek-R1 [guo2025deepseek], integrate chain-of-thought and reinforcement learning to enhance high-level inference. Building on these advances, GPT-4O [hurst2024gpt] and Gemini 2.5 [comanici2025gemini] combine perception and reasoning to support complex, agentic decision-making. Collectively, these milestones progressively enable hierarchical spatial intelligence in MLLMs, motivating structured benchmarks and evaluation frameworks across low-level perception, intermediate reasoning, and high-level agentic competence.
@@
Benchmarks for spatial abilities in MLLMs have evolved alongside the models themselves. Early efforts, such as BLINK [fu2024blink], SpatialEval [wang2024spatial], and 3DSR-Bench [ma20243dsrbench], focused on evaluating spatial understanding tasks in single images, including distance estimation, relational question answering, and spatial captions. As MLLMs increasingly support multi-frame and video inputs, benchmarks such as VSI-Bench [yang2025thinking] and MMSI-Bench [yang2025mmsi] have emerged to evaluate spatial reasoning across multiple views and dynamic scenes. To further enrich task diversity and coverage, Omnispatial [jia2025omnispatial], SITE [wang2025site], and IR3D-Bench [liu2025ir3d] extend benchmarks to geometry puzzles, dynamic reasoning, and inverse rendering tasks. Built upon prior efforts, our SpatialTree benchmark systematically organizes spatial abilities into a hierarchical framework, providing the first thorough evaluation across different capabilities.
@@ 3.4 Agentic Competence
where $O_{t}\in\mathcal{O}$ is the current multi-modal observation, $S_{t}\in\mathcal{S}$ the internal latent state (e.g., goal, plan, or belief), $A_{t}\in\mathcal{A}$ the chosen action, and $M_{t}\in\mathcal{M}$ the updated memory representation. MLLMs are expected to output interactive actions executable across 3D environments and embodiments, such as games, simulators, and the physical world. Unlike Vision-Language Action Models (VLAs) decording the low-level control signals in robotics [intelligence2025pi_], MLLMs take the language as the only interface to link with environments like GUI Agents [UI-TARS].
@@ 4.1 Data Curation
Data Annotation Framework. As shown in Fig. 3, the data engines are organized hierarchically. At the perception level, we employ a set of expert perception models, including DepthAnything3 [yang2024depth], SpatialTracker [xiao2025spatialtrackerv2, xiao2024spatialtracker], GeoCalib [veicht2024geocalib], and OrientAnything [orient_anything], to extract intermediate perceptual representations, such as depth, correspondences, tracking results, and gravity direction. These representations are then further processed using our designed QA templates and rephrased by LLMs to construct the target QAs. For L2, we leverage 3D reconstruction pipelines (Sec. C.2) to generate BEV maps from videos, which are subsequently captioned and transformed into QAs for cognitive mapping and memory retrieval with multimodal LLMs. For other understanding tasks, such as affordance, we build on partially annotated datasets [luo2022learning] and augment them by randomly constructing multiple-choice candidates through various visual prompts and by reformulating questions with more abstract descriptions. For L3, we build upon annotated reasoning-related QAs and introduce structured thinking templates, enabling an LLM-based rephraser to augment the original QAs with explicit Chain-of-Thought (CoT) reasoning. For agentic tasks, we curate Internet data covering manipulation and navigation across diverse embodiments (e.g., human hands, robotic grippers, and game characters). We further process the data with our Action-Extraction Pipeline (Sec. C.2), converting embodiment-specific actions into unified key–mouse action sequences suitable for MLLMs. In parallel, we introduce a human annotation pipeline to annotate videos with executable multi-option action sequences.
Data Resources. Our SpatialTree-Bench is constructed by systematically reorganizing numerous recent datasets (detailed in the Appendix) [yang2025thinking, yang2025mmsi, zhu2024llava, wang2024spatial, yin2025spatial, lin2025towards, jia2025omnispatial, yang2025embodiedbench, wang2025spatialviz, xu2025multi, ma20243dsrbench] to address their scattered capability coverage and over-reliance on simple questions. We first map each question to our SpatialTree framework and then enhance the evaluation protocol; for instance, complex reasoning tasks from CameraBench and MMSI-Bench are converted to a hybrid multi-option + LLM-as-a-Judge format for a finer-grained assessment. To fill the remaining capability gaps, we introduce SpatialPlus, a new dataset targeting underrepresented abilities (e.g., L1 Orientation, L1 Shape, L2 Spatial Caption) with a primary emphasis on L4 Agentic Competence. We leverage our proprietary SpatialEngine to automatically create annotations from a diverse array of video sources, including 3D reconstruction datasets, in-game footage [ju2024miradata], egocentric manipulation videos [egodex], and robotics data [khazatsky2024droid]. More implementation details are discussed in Appendix.
@@ 5.1 Models and Metrics
Benchmarked Models. We categorize the evaluated MLLMs into three groups: (1) Thinking Models, i.e., models augmented with explicit reasoning or chain-of-thought generation mechanisms (reasoning-augmented), including Gemini 3 Pro [gemini3pro2026], Gemini 3 Flash [gemini3flash2026], Gemini 2.5 Flash, Gemini 2.5 Pro [comanici2025gemini], GLM-4.5V [hong2025glm], Seed1.5VL [guo2025seed1] and Seed1.8 [seedseed1]; (2) Non-Thinking Models, which do not explicitly optimize for reasoning-style generation, such as Gemini 2.5-Pro-Nonthink, Gemini 2.5-Flash-Nonthinking, and GPT-4o [hurst2024gpt]; and (3) Open-Source Models, including Qwen2.5-VL [bai2025qwen2], Qwen3-VL [yang2025qwen3], and Kimi-VL [team2025kimi], representing recent community-driven multimodal advances. This diverse selection enables comprehensive comparisons across reasoning and non-reasoning paradigms, proprietary and open-source ecosystems, and model scales ranging from 32B to 72B parameters—providing a holistic overview of the current MLLM landscape. A full list of evaluated models is shown in Tab. 5.2.
@@ 5.2 Overall Performance
Seed1.5vl 6 41.3 39.2 36.6 24.6 44.7 44.8 46.5 38.5 36.6 47.0 28.4 25.9
[…]
GLM4.5V 9 36.0 35.3 24.0 32.5 34.5 43.7 34.5 34.1 33.8 41.0 26.8 49.7
Gemini2.5-Pro 4 50.1 47.8 32.6 44.4 61.6 47.9 50.5 61.5 47.6 58.2 28.3 63.3
Gemini2.5-Flash 7 41.1 35.6 29.3 40.5 56.2 44.6 43.1 51.4 36.4 49.1 26.9 46.6
@@
We first present the overall performance of all benchmarked models on our proposed SpatialTree-Bench, with detailed results summarized in Tab. 5.2. In our benchmark, Gemini 3 Flash achieves the best results (57.8) and Qwen3VL-235B get 40.0 leads the benchmark among open-source models.
@@ 6.1 Probing Cross-Ability Transfer via SFT
General Data Mixture. To construct the general visual-instruction data, we follow the VST [vst_yang2025] data mixing recipe and combine multimodal datasets from LLaVA-Video [zhang2024llavanext-video], LLaVA-NeXT-Interleave [li2024llava_next_interleave], and LLaVA-OneVision [li2024llava_onevision], covering single-image, multi-image, video, and 3D tasks. We then use SpatialEngine to generate ability-specific instruction data, mixed with the general data in a 1:3 ratio specifically for each. To isolate the gains from general data, we use a baseline fine-tuned only on the general data with the same token consumption.
Targeted SFT Data. For L1-Geo.Dist, we generate approximately 0.25M distance-relevant QA samples from SUNRGBD [song2015sun], Hypersim [Hypersim], and Matterport3D [Matterport3D]. The training data is further augmented with visual prompts and multi-scale transformations to enhance distance reasoning. For L1-Relat.Corr, we generate matching data following VST [vst_yang2025], sampling 0.25M examples. Similarly, for L1-Geo.Size, we generate 0.25M samples from 3D bounding-box annotated datasets, including SUNRGBD, Hypersim and ArkitScenes [baruch2021arkitscenes].
@@ 6.2 Reinforcement Learning
Building on the findings from Supervised Fine-Tuning (SFT), we further investigate the potential of Reinforcement Learning with Verifiable Rewards (RLVR) to scale spatial abilities. Specifically, we employ Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] to align the model’s policy with the hierarchical nature of the SpatialTree. Our experiments reveal that a uniform RL strategy is insufficient, conversely a hierarchy-aware reward mechanism unlocks significant performance gains.
@@ B Evaluation Metrics Details
To assess agentic competence, models are deployed in interactive simulated environments, such as those provided by EmbodiedBench [yang2025embodiedbench]. We evaluate navigation and manipulation tasks along multiple dimensions: success rate in completing the target goal, relative translation accuracy, and directional alignment. For each action step, a combined metric is computed using relative distance and cosine similarity of movement vectors, producing a normalized score in $[0,1]$. Aggregating scores over all steps yields a comprehensive measure of an agent’s ability to plan and execute actions in long-horizon, embodied tasks.
@@ C SpatialPlus: Complementary Data Annotations for SpatialTree
Figure B: Orientation Annotations. The left side is the gravity field estimated from GeoCalib [veicht2024geocalib], while the right side is from OrientAnything.
@@ C.1 Orientations (L1)
The Orientation capability, a fundamental yet under-explored area, involves estimating both gravity direction and 3D object orientation. To generate annotations, we leveraged Geocalib [veicht2024geocalib] for gravity vector estimation and OrientAnything [orient_anything] for object poses. We applied these tools to datasets suited for each task: for gravity, we annotated 500 images sampled from the diverse, drone-captured TartanAir [wang2020tartanair] dataset; for object orientation, we utilized the object-centric Co3dv2 [reizenstein21co3d] dataset (Seen in Fig. B). For gravity, the goal is to estimate the camera’s orientation relative to the gravity vector, typically represented by the pitch and roll angles. Formally, let the gravity vector in the world frame be: